# ToDRE: Visual Token Pruning via Diversity and Task Awareness for Efficient Large Vision-Language Models

### v1

@@ C.2 Layers in Relevance-driven Visual Token Reduction
LLaVA-NeXT-7B [37] 77:04 1519.6 72.0 64.2 87.7 100.0%
$L/2$ 79:03 1503.9 70.6 64.0 87.7 99.2%
$85L/$ 8 79:00 1521.5 71.9 64.0 87.7 99.9%
$86L/$ 8 79:01 1518.4 71.8 64.2 87.6 99.9%
$87L/$ 8 78:38 1513.7 71.9 64.3 87.6 100.0%
$L$ 77:03 1519.3 72.0 64.2 87.6 100.0%
$L/2\sim L$ 155:01 1028.7 67.5 43.1 60.3 74.3%
$L/2+5L/8+6L/8+7L/8+L$ 74:13 1495.0 70.5 63.7 87.8 98.9%
$\mathbf{L/2+5L/8+6L/8+7L/8}$ (Ours) 72:46 1497.5 70.5 63.7 87.7 98.9%
Table 6: Ablation study on selected layers in relevance-driven visual token reduction. $L$ denotes the total number of decoder layers in the LLM. The bottom row corresponds to the default setting used in the main paper.
We conduct an ablation study to investigate the optimal candidate layer selection when applying relevance-driven visual token reduction strategy (Section 4.2) during LLM prefilling. As shown in Table 6, applying the adaptive strategy at a single layer (i.e., from ‘$L/2$’ to ‘$L$’ rows) is insufficient, since the attention ratios $\alpha_{v\to t}$ and $\alpha_{t\to v}$ (Equation 7)-which represent the extent of cross-modal interaction-exhibit fluctuations across layers. Thus, relying on a single-layer measurement yields longer inference time and cannot consistently capture the necessary conditions for effective token pruning. Conversely, implementing the adaptive strategy at every deep layer (i.e., row ‘$L/2\sim L$’) introduces excessive computational overhead, undermining and even reversing the efficiency benefits provided by token pruning.
Instead, we observe that applying the proposed strategy at interleaved layers (i.e., the last two rows) achieves a favorable balance between computational cost and model performance. Empirical observations further indicate that including the final layer does not yield additional performance improvements but instead adds unnecessary computational expense. Thus, the layer selection strategy described in Section 4.2 offers the optimal trade-off between performance and efficiency.
Appendix D Negligible Cross-Attention to Visual Tokens during LLM Decoding
@@
Figure 5: The output token’s attention towards different input token types across LLM layers in decoding stage. The visualization is based on four benchmarks: MME [14], SQA [41], GQA [21], and POPE [31], with 100 samples randomly selected from each. Attention ratios are averaged over all samples within each benchmark.
@@
Building on the clear information migration phenomenon observed during the LLM prefilling stage (Figure 2), we further examine the effect of pruning all visual tokens during decoding stage. As shown in Figure 5, in the shallow layers, attention from output tokens to system prompt tokens increases sharply, while attention directed towards visual tokens drops significantly. Moreover, in the middle and deeper layers, the output tokens consistently exhibit high attention towards system prompt tokens and text tokens, whereas attention to visual tokens remains negligible (less than 5%). These findings further validate the validity of our relevance-driven visual token reduction strategy.
Appendix E More Case Studies
E.1 Free-Form Question Answering with Long Responses
@@
Figure 6: Qualitative comparison of free-form video-grounded QA on the Video Detail Caption benchmark [9]. Green text highlights correctly identified events and objects; red text indicates incorrect predictions; yellow text marks missing but essential information.
@@
We present a qualitative comparisons of free-form question answering with long responses on the Video Detail Caption benchmark [9]. As shown in Figure 6, our method accurately identifies both the event and activity depicted in the video. In contrast, FastV [10] produces a vague description of the action and omits key objects, while FasterVLM [61] generates a generic caption (“a throwing motion”) and incorrectly identifies the main object. These comparisons highlight the superior descriptive precision of our approach in capturing fine-grained visual details.
@@
Figure 7: Supplementary visualizations of attention- and diversity-driven token compression. The visualization is based on seven benchmarks: MME [14], SQA [41], GQA [21], POPE [31], MMBench and MMBench-CN [40], VQAv2 [17]. Best viewed when zoomed in.
E.2 Attention-Driven Token Pruning vs. Diversity-Driven Token Retention
@@
We provide additional visualizations of attention-based token reduction and diversity-driven token retention across various image understanding benchmarks, including MME [14], SQA [41], GQA [21], POPE [31], MMBench and MMBench-CN [40], VQAv2 [17]. As shown in Figure 7, attention-based token retention tends to produce more concentrated token distributions, focusing on a limited subset of high-attention regions. In contrast, diversity-based retention results in a more dispersed selection of tokens, covering broader spatial and semantic regions. This broader coverage enables the model to better handle a wider array of open-ended questions.
Appendix F Limitation
Compared to training-based visual token compression methods, training-free approaches can be seamlessly integrated into various LVLMs. However, they often inevitably introduce additional computational overhead, such as computing pairwise similarity matrices or running statistical algorithms, which can partially offset the gains in inference efficiency.

### v2

@@ 3.3 Layers in Relevance-driven Visual Token Reduction
[…]
$L$ 78:51 1519.6 71.8 64.2 87.6 99.9%
$L/2\sim L$ 39:20 1527.8 65.4 39.0 87.1 87.9%
$L/2+5L/8+6L/8+7L/8$ 65:22 1527.8 65.4 39.1 87.2 87.9%
$L/2$ 64:19 1527.8 65.4 39.0 87.1 87.9%
$5L/8$ 58:14 1528.6 71.8 54.3 87.5 96.2%
$6L/8$ 68:54 1527.3 71.8 59.8 87.6 98.3%
$\mathbf{7L/8}$ (Ours) 70:15 1522.7 71.9 64.3 87.6 100.0%
Table 3: Ablation study on selected layers in relevance-driven visual token reduction. $L$ denotes the total number of decoder layers in the LLM. The bolded row corresponds to the default setting used in the main paper.
We conduct an ablation study to investigate the optimal candidate layer selection when applying relevance-driven visual token reduction strategy during LLM prefilling. As shown in Table˜3, applying the adaptive reduction at earlier layers (e.g., starting from ‘$L/2$’) is suboptimal, as the attention ratios $\alpha_{v\to t}$ and $\alpha_{t\to v}$—which characterize the degree of cross-modal interaction—fluctuate considerably in the early layers. Early pruning thus prematurely interrupts the ongoing alignment process between modalities.
In contrast, applying the proposed strategy at later decoder layers (i.e., the last three rows) yields a favorable trade-off between efficiency and performance. Although pruning at $5L/8$ or $6L/8$ further reduces the inference time, both settings incur a noticeable drop in average performance (around 2–4% compared to the full model). By contrast, applying the reduction at $7L/8$ restores the performance to the baseline level while maintaining nearly the same computational efficiency. Therefore, the layer selection strategy described in the main paper—starting from 7L/87L/8—achieves the best balance between accuracy and efficiency.
4 Negligible Cross-Attention to Visual Tokens during LLM Decoding
@@
Figure 1: Output token’s attention toward different input token types across LLM layers during decoding. Results are averaged over 100 samples per benchmark.
@@
Building on the clear information migration phenomenon observed during the LLM prefilling stage, we further examine the effect of pruning all visual tokens during decoding stage. As shown in Figure˜1, in the shallow layers, attention from output tokens to system prompt tokens increases sharply, while attention directed towards visual tokens drops significantly. Moreover, in the middle and deeper layers, the output tokens consistently exhibit high attention towards system prompt tokens and text tokens, whereas attention to visual tokens remains negligible (less than 5%). These findings further validate the effectiveness of our relevance-driven visual token reduction strategy.
5 More Case Studies
5.1 Free-Form Question Answering with Long Responses
@@
Figure 2: Qualitative comparison of free-form video-grounded QA on the Video Detail Caption benchmark [chai2025videodetailcaption]. Green text highlights correctly identified events and objects; red text indicates incorrect predictions; yellow text marks missing but essential information.
@@
We present qualitative comparisons of free-form question answering with long responses on the Video Detail Caption benchmark [chai2025videodetailcaption]. As shown in Figure˜2, our method accurately identifies both the event and activity depicted in the video. In contrast, FastV [chen2024fastv] produces a vague description of the action and omits key objects, while FasterVLM [zhang2024fastervlm] generates a generic caption (“a throwing motion”) and incorrectly identifies the main object. These comparisons highlight the superior descriptive precision of our approach in capturing fine-grained visual details.
@@
Figure 3: Supplementary visualizations comparing attention-driven and ToDRE-based token compression. The visualization is based on seven benchmarks: MME [Fu2023mme], SQA [lu2022sqa], GQA [hudson2019gqa], POPE [li2023pope], MMBench and MMBench-CN [Liu2023mmbench], VQAv2 [goyal2017vqav2]. Best viewed when zoomed in.
5.2 Attention-Driven Token Pruning vs. ToDRE Token Retention
@@
We provide additional visualizations of attention-based token reduction and ToDRE token retention across various image understanding benchmarks, including MME [Fu2023mme], SQA [lu2022sqa], GQA [hudson2019gqa], POPE [li2023pope], MMBench and MMBench-CN [Liu2023mmbench], VQAv2 [goyal2017vqav2]. As shown in Figure˜3, attention-based token retention tends to produce more concentrated token distributions, focusing on a limited subset of high-attention regions. In contrast, ToDRE retention results in a more dispersed selection of tokens, covering broader spatial and semantic regions. This broader coverage enables the model to better handle a wider array of open-ended questions.